OpenStudy (anonymous):

find sec and csc when tan=3?

OpenStudy (lgbasallote):

if tan = 3 then that means opposite side is 3 and adjacent side is 1. make sense?

OpenStudy (anonymous):

i think so it's one because on the unit circle the radius is one i guess? that's how i'm thinking about it

OpenStudy (lgbasallote):

no.. that's actually because \[\tan = 3 = \frac 31 = \frac{opposite}{adjacent}\] so opposite side is 3 and adjacent side is 1

OpenStudy (lgbasallote):

does it make sense now?

OpenStudy (anonymous):

yeah okay that makes more sense

OpenStudy (lgbasallote):

good so that means hypotenuse is \[\huge \sqrt{3^2 + 1^2}\implies \sqrt{9 + 1} \implies \sqrt {10}\] right?

OpenStudy (anonymous):

yeah that makes comeplete sense! okay so in another question i have to find tan and cot when sec=4 so does that mean....

OpenStudy (lgbasallote):

\[\sec = 4 = \frac 41 = \frac{hypotenuse}{adjacent}\] so that means hypotenuse = 4 and adjacent = 1 right?

OpenStudy (anonymous):

the hypotenuse is 4 and the adjacent angle is 1?

OpenStudy (lgbasallote):

now....use pythagorean theorem \[\huge a^2 + b^2 = c^2\] \[\huge \implies a^2 = c^2 - b^2\] \[\huge \implies a = \sqrt{c^2 - b^2}\] got it?

OpenStudy (anonymous):

yeah okay i think i getting this! okay last thing, what if i get something wierd like cos=0.8?

OpenStudy (lgbasallote):

0.8 means 8/10

OpenStudy (lgbasallote):

you turn them into fractions

OpenStudy (anonymous):

okay so whenever i get a number i think of that number over what gives me that right?

OpenStudy (lgbasallote):

hmm many word plays but it sounds about right

OpenStudy (anonymous):

okay cool thank you so much!

OpenStudy (lgbasallote):

welcome